OpenStudy (anonymous):

One number is greater than half of a another number by 15. What are the 2 numbers if their sum is 48? I need the equations and answer pleaseee :))

OpenStudy (anonymous):

didn't you just ask this question?

OpenStudy (anonymous):

m+n=48?

OpenStudy (anonymous):

m/2+15=n?

OpenStudy (anonymous):

n is greater than m/2 by fifteen

OpenStudy (anonymous):

substitute m/2+15 for n in the original equation m+n=48